OpenStudy (anonymous):

Evaluate 6x ÷ 3y - 4, if x = 20 and y = 2.

OpenStudy (destinymasha):

6 x 20 divided by 3 x 2 -4

OpenStudy (anonymous):

its 16

OpenStudy (anonymous):

(6 * 20) = 120 (3 * 2) = 6 So: 120 / 6 = 20 20 - 4 = 16 16 is your answer.

OpenStudy (destinymasha):

sorry guys..

OpenStudy (destinymasha):

its 16